OpenStudy (anonymous):

would this be a true statement? \[u=\sqrt{6}{x}\] therefore... \[u^2=\sqrt{3}{x}\]

OpenStudy (anonymous):

i meant the 6th root

OpenStudy (anonymous):

typo

OpenStudy (anonymous):

sixth root of x and 3rd root of x

OpenStudy (unklerhaukus):

±

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

thats right

OpenStudy (anonymous):

yeah i guess \[\frac 1 2 *\frac 1 3 \]

OpenStudy (anonymous):

no

OpenStudy (unklerhaukus):

\[u=\pm\sqrt[6]{x}\]\[u^2=\sqrt[3]{x}\]

OpenStudy (anonymous):

i mean \[x^{\frac 1 2 *\frac 1 3}\]

OpenStudy (anonymous):

just notice \[\sqrt[a]{x}=x^{1/a}\] and \[(x^b)^c=x^{bc}\]

OpenStudy (anonymous):

i agree

OpenStudy (anonymous):

\[u=x^{\frac 1 2 *\frac 1 3}\] correct?

OpenStudy (anonymous):

why not?

OpenStudy (anonymous):

i my notation wrong or the idea all together

OpenStudy (anonymous):

sorry thats right

OpenStudy (anonymous):

well what then?

OpenStudy (anonymous):

nothing. \[u^2=\sqrt{3}{x}\] (cube root) didn't look right to me at first but then it made sense...thanks!